# Latent Schrödinger Bridge Diffusion Model for Generative Learning

### v1

@@ C.2 Statistical Error
$\mathbb{E}_{\mathbf{z}}\left\|\mathbf{s}(t,\widehat{\mathbf{x}}+\sigma\sqrt{1-t}\mathbf{z})+\frac{\mathbf{z}}{\sigma\sqrt{1-t}}\right\|^{2}$ $\leq{2}\left(K^{2}+\frac{\mathbb{E}\|\mathbf{z}\|^{2}}{\sigma^{2}(1-t)}\right)$
$=2\left(K^{2}+\frac{d}{\sigma^{2}(1-t)}\right).$
@@
$\ell_{\mathbf{s}}(\widehat{\mathbf{x}})\leq C_{T}:=2\left(K^{2}-\frac{d^{*}\log{(1-T)}}{\sigma^{2}T}\right).$
@@
Now we derive the Lipschitz continuity for $\ell_{\mathbf{s}}(\widehat{\mathbf{x}})$. We have
@@
$\leq\frac{1}{T}\int_{0}^{T}\left(\mathbb{E}_{\mathbf{z}}\|\mathbf{s}_{1}-\mathbf{s}_{2}\|^{2}\right)^{\frac{1}{2}}\left(6K^{2}+\frac{12d^{*}}{\sigma^{2}(1-t)}\right)^{\frac{1}{2}}\mathrm{d}t$
$\leq\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}\left\|\mathbf{s}_{1}-\mathbf{s}_{2}\right\|^{2}\mathrm{d}t\right)^{\frac{1}{2}}\left(6K^{2}+\frac{12d^{*}}{T}\int_{0}^{T}\frac{1}{\sigma^{2}(1-t)}\mathrm{d}t\right)^{\frac{1}{2}}$
$\leq\sqrt{6K^{2}-\frac{12d^{*}\log(1-T)}{T\sigma^{2}}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}\|\mathbf{s}_{1}-\mathbf{s}_{2}\|^{2}\mathrm{d}t\right)^{\frac{1}{2}}$
$\leq\sqrt{6K^{2}-\frac{12d^{*}\log(1-T)}{T\sigma^{2}}}\|\mathbf{s}_{1}-\mathbf{s}_{2}\|_{L^{\infty}([0,T]\times\mathbb{R}^{d^{*}})}$
$=\sqrt{6K^{2}-\frac{12d^{*}\log(1-T)}{T\sigma^{2}}}\|\mathbf{s}_{1}-\mathbf{s}_{2}\|_{L^{\infty}([0,T]\times[-r,r]^{d^{*}})}.$
For convenience, we let
$B_{T}:=\sqrt{6K^{2}-\frac{12d^{*}\log(1-T)}{T\sigma^{2}}}.$
@@
$\log\mathcal{N}\left(\text{NN},\delta,\|\cdot\|_{L^{\infty}([0,T]\times\mathbb{R}^{d^{*}}}\right)$ $=\log\mathcal{N}\left(\text{NN},\delta,\|\cdot\|_{L^{\infty}([0,T]\times[-r,r]^{d^{*}})}\right)$
$\lesssim JL\log{\left(\frac{LMr\kappa}{\delta}\right)}$
$=\widetilde{\mathcal{O}}\left(\frac{\xi^{d}\epsilon^{-(d^{*}+1)}}{(1-T)^{3}}\right),$
where we choose $r=\mathcal{O}\left(\sqrt{\log\frac{d^{*}}{\epsilon(1-T)}}\right)$ larger than 1. The details of this derivation can be found in [CJLZ22, Lemma 5.3].
Upper bound for statistical error.
Let $\ell(\mathbf{s},\widehat{\mathbf{x}})=\ell_{\mathbf{s}}(\widehat{\mathbf{x}})-\ell_{\mathbf{s}^{*}}(\widehat{\mathbf{x}})$ and $\widehat{\mathcal{X}}^{\prime}=\{\widehat{\mathbf{x}}_{1}^{\prime},\cdots,\widehat{\mathbf{x}}_{n}^{\prime}\}$ be an independent copy of $\widehat{\mathcal{X}}$, then
$|\ell(\mathbf{s}_{1},\widehat{\mathbf{x}})-\ell(\mathbf{s}_{2},\widehat{\mathbf{x})}|=|\ell_{\mathbf{s}_{1}}(\widehat{\mathbf{x}})-\ell_{\mathbf{s}_{2}}(\widehat{\mathbf{x}})|\leq B_{T}\|\mathbf{s}_{1}-\mathbf{s}_{2}\|_{L^{\infty}([0,T]\times\mathbb{R}^{d^{*}})}.$
@@
where $G(\widehat{\mathbf{s}},\widehat{\mathbf{x}})=\mathbb{E}_{\widehat{\mathcal{X}}^{\prime}}\left[\ell(\widehat{\mathbf{s}},\widehat{\mathbf{x}}^{\prime})-2\ell(\mathbf{s},\widehat{\mathbf{x}})\right].$
Let $\mathcal{C}_{\delta}$ be the $\delta$-covering of $\mathrm{NN}$ with minimum cardinality $\mathcal{N}:=\mathcal{N}(\mathrm{NN},\delta,\|\cdot\|_{L^{\infty}([0,T]\times\mathbb{R}^{d^{*}})})$, then for any $\mathbf{s}\in\mathrm{NN}$, there exists a $\mathbf{s}_{\delta}$ such that
$|\ell(\mathbf{s},\widehat{\mathbf{x}})-\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}})|\leq B_{T}\|\mathbf{s}-\mathbf{s}_{\delta}\|_{L^{\infty}([0,T]\times\mathbb{R}^{d^{*}})}\leq B_{T}\delta.$
@@
$G(\widehat{\mathbf{s}},\widehat{\mathbf{x}})\leq G(\mathbf{s}_{\delta},\widehat{\mathbf{x}})+3B_{T}\delta.$
Since $|\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})|\leq 2C_{T}$, we have $|\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})-\mathbb{E}\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})|\leq 4C_{T}$. Let $V^{2}=\mathrm{Var}[\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})]$, then
@@
$\leq B_{T}^{2}\mathbb{E}_{\widehat{\mathcal{X}}}\left[\ell_{\mathbf{s}_{\delta}}(\widehat{\mathbf{x}}_{i})-\ell_{\mathbf{s}^{*}}(\widehat{\mathbf{x}}_{i})\right]$
$=B_{T}^{2}\mathbb{E}_{\widehat{\mathcal{X}}}[\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})].$
@@
$\mathbb{E}_{\widehat{\mathcal{X}}}[\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})]\geq\frac{V^{2}}{B_{T}^{2}}.$
@@
$\leq$ $\leavevmode\nobreak\ \mathbb{P}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\mathbb{E}_{\widehat{\mathcal{X}}}\left[\frac{1}{n}\sum_{i=1}^{n}\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})\right]-\frac{1}{n}\sum_{i=1}^{n}\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})>\frac{t}{2}+\frac{V^{2}}{2B_{T}^{2}}\right)$
$\leq$ $\leavevmode\nobreak\ \exp\left(-\frac{nu^{2}}{2V^{2}+\frac{8uC_{T}}{3}}\right)$
$\leq$ $\exp\left(-\frac{nt}{8B_{T}^{2}+\frac{16C_{T}}{3}}\right),$
where $u=\frac{t}{2}+\frac{V^{2}}{2B_{T}^{2}}$, and we use $u\geq\frac{t}{2}$ and $V^{2}\leq 2B_{T}^{2}u$. Hence, for any $t>3B_{T}\delta$, we have
$\mathbb{P}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\frac{1}{n}\sum_{i=1}^{n}G(\widehat{\mathbf{s}},\widehat{\mathbf{x}}_{i})>t\right]$ $\leq\mathbb{P}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\mathop{\mathrm{sup}}_{\mathbf{s}\in\mathrm{NN}}\frac{1}{n}\sum_{i=1}^{n}G(\mathbf{s},\widehat{\mathbf{x}}_{i})>t\right]$
$\leq\mathbb{P}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\mathop{\mathrm{max}}_{\mathbf{s}_{\delta}\in\mathcal{C}_{\delta}}\frac{1}{n}\sum_{i=1}^{n}G(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})>t-3B_{T}\delta\right]$
$\leq\mathcal{N}\mathop{\mathrm{max}}_{\mathbf{s}_{\delta}\in\mathcal{C}_{\delta}}\mathbb{P}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\frac{1}{n}\sum_{i=1}^{n}G(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})>t-3B_{T}\delta\right]$
$\leq\mathcal{N}\exp\left(-\frac{n(t-3B_{T}\delta)}{8B_{T}^{2}+\frac{16C_{T}}{3}}\right).$
Setting $a=3B_{T}\delta+\frac{\left(8B_{T}^{2}+\frac{16C_{T}}{3}\right)\log\mathcal{N}}{n}$ and $\delta=\frac{1}{n}$, then we obtain
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\frac{1}{n}\sum_{i=1}^{n}G(\widehat{\mathbf{s}},\widehat{\mathbf{x}}_{i})\right]$ $\leq a+\mathcal{N}\int_{a}^{\infty}\exp\left(-\frac{n(t-3B_{T}\delta)}{8B_{T}^{2}+\frac{16C_{T}}{3}}\right)\mathrm{d}t$
$\leq\frac{\left(8B_{T}^{2}+\frac{16C_{T}}{3}\right)\left(\log\mathcal{N}+1\right)+3B_{T}}{n}$
$=\widetilde{\mathcal{O}}\left(\frac{1}{n}\cdot\frac{\xi^{d^{*}}\epsilon^{-(d^{*}+1)}}{(1-T)^{5}}\right).$
@@
$\ell_{\widehat{\mathbf{s}}}^{\mathrm{trunc}}(\widehat{\mathbf{x}}_{i})=\mathbb{E}_{\mathbf{z}}\left(\frac{1}{T}\int_{0}^{T}\left\|\widehat{\mathbf{s}}(t,\widehat{\mathbf{x}}_{i}+\sigma\sqrt{1-t}\mathbf{z})+\frac{\mathbf{z}}{\sigma\sqrt{1-t}}\right\|^{2}\mathrm{d}t\cdot\mathbf{I}_{\{\|\mathbf{z}\|_{\infty}\leq r_{0}\}}\right),$
@@
^$\widehat{\ell}_{\widehat{\mathbf{s}}}^{\mathrm{trunc}}(\widehat{\mathbf{x}}_{i})=\frac{1}{m}\sum_{j=1}^{m}\left\|\widehat{\mathbf{s}}(t_{j},\widehat{\mathbf{x}}_{i}+\sigma\sqrt{1-t_{j}}\mathbf{z}_{j})+\frac{\mathbf{z}_{j}}{\sigma\sqrt{1-t_{j}}}\right\|^{2}\mathbf{I}_{\{\|\mathbf{z}_{j}\|_{\infty}\leq r_{0}\}}.$
@@
$(A)$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\mathbf{z}}\left(\frac{1}{T}\int_{0}^{T}\left\|\widehat{\mathbf{s}}(t,\widehat{\mathbf{x}}_{i}+\sigma\sqrt{1-t}\mathbf{z})+\frac{\mathbf{z}}{\sigma\sqrt{1-t}}\right\|^{2}\mathrm{d}t\cdot\mathbf{I}_{\{\|\mathbf{z}\|_{\infty}>r_{0}\}}\right)$
$\leq 2K^{2}\mathbb{P}(\|\mathbf{z}\|_{\infty}>r_{0})+\frac{2}{T}\mathbb{E}\left(\|\mathbf{z}\|^{2}\mathbf{I}_{\{\|\mathbf{z}\|_{\infty}>r_{0}\}}\right)\int_{0}^{T}\frac{1}{\sigma^{2}(1-t)}\mathrm{d}t$
$\leq 2K^{2}\mathbb{P}(\|\mathbf{z}\|_{\infty}>r_{0})-\frac{2\log{(1-T})}{\sigma^{2}T}[\mathbb{E}(\|\mathbf{z}\|^{4})]^{\frac{1}{2}}\mathbb{P}({\|\mathbf{z}\|_{\infty}>r_{0}})^{\frac{1}{2}}$
$\leq\left(2K^{2}-\frac{2\log{(1-T)}\sqrt{d^{*}(d^{*}+2)}}{\sigma^{2}T}\right)\mathbb{P}({\|\mathbf{z}\|_{\infty}>r_{0}})^{\frac{1}{2}}$
$\leq\left(2K^{2}-\frac{2\log{(1-T)}\sqrt{d^{*}(d^{*}+2)}}{\sigma^{2}T}\right)\sqrt{2d^{*}}\exp\left(-\frac{r_{0}^{2}}{4}\right).$
Therefore,
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\frac{1}{n}\sum_{i=1}^{n}\left(\ell_{\widehat{\mathbf{s}}}(\widehat{\mathbf{x}}_{i})-\ell_{\widehat{\mathbf{s}}}^{\mathrm{trunc}}(\widehat{\mathbf{x}}_{i})\right)\right)$
$\leq\left(2K^{2}-\frac{2\log{(1-T)}\sqrt{d^{*}(d^{*}+2)}}{\sigma^{2}T}\right)\sqrt{2d^{*}}\exp\left(-\frac{r_{0}^{2}}{4}\right).$
Let $h_{\mathbf{s}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})=\|\mathbf{s}(\widehat{\mathbf{x}}_{i}+\sigma\sqrt{1-t}\mathbf{z})+\frac{\mathbf{z}}{\sigma\sqrt{1-t}}\|^{2}\mathbf{I}_{\{\|\mathbf{z}\|_{\infty}\leq r_{0}\}}$, then
$h_{\mathbf{s}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})\leq E_{T}(r_{0}):=2K^{2}+\frac{d^{*}r_{0}^{2}}{\sigma^{2}(1-T)}.$
For any $\mathbf{s}\in\mathrm{NN}$, there exists a $\mathbf{s}_{\delta}\in\mathcal{C}_{\delta}$ such that
$|h_{\mathbf{s}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})-h_{\mathbf{s}_{\delta}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})|$ $\leq\delta\left\|\mathbf{s}+\mathbf{s}_{\delta}+\frac{2\mathbf{z}}{\sigma\sqrt{1-t}}\right\|\mathbf{I}_{\{\|\mathbf{z}\|_{\infty}\leq r_{0}\}}$
$\leq\left(2K+\frac{2\sqrt{d^{*}}r_{0}}{\sigma\sqrt{1-T}}\right)\delta.$

### v2

@@ Appendix C Statistical Error
$\mathbb{E}_{\mathbf{z}}\left\|\mathbf{s}(1-t,\widehat{\mathbf{x}}+\sigma\sqrt{1-t}\mathbf{z})+\frac{\mathbf{z}}{\sigma\sqrt{1-t}}\right\|^{2}\lesssim\frac{\log\mathcal{R}}{\sigma^{2}(1-t)}+\frac{\mathbb{E}\|\mathbf{z}\|^{2}}{\sigma^{2}(1-t)}\lesssim\frac{\log\mathcal{R}}{1-t}.$
@@
$\ell_{\mathbf{s}}(\widehat{\mathbf{x}})\lesssim\frac{\log^{2}\mathcal{R}}{1-{\mathcal{R}}^{-C_{T}}}\lesssim\log^{2}\mathcal{R}$
for sufficiently large $\mathcal{R}$.
@@
Now we derive the Lipschitz continuity for $\ell_{\mathbf{s}}(\widehat{\mathbf{x}})$. Note that we restrict ReLU neural networks into class $\mathcal{C}$. For any $\mathbf{s}_{1}$, $\mathbf{s}_{2}\in\mathcal{C}$, by the construction structure of $\mathbf{s}_{1}$, $\mathbf{s}_{2}$, we have
@@
$\lesssim\frac{1}{T}\int_{0}^{T}\left(\mathbb{E}_{\mathbf{z}}\|\mathbf{s}_{1}-\mathbf{s}_{2}\|^{2}\right)^{\frac{1}{2}}\left(\frac{\log\mathcal{R}}{\sigma^{2}(1-t)}+\frac{d^{*}}{\sigma^{2}(1-t)}\right)^{\frac{1}{2}}\mathrm{d}t$
$\lesssim\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}\left\|\mathbf{s}_{1}-\mathbf{s}_{2}\right\|^{2}\mathrm{d}t\right)^{\frac{1}{2}}\left(\log\mathcal{R}\cdot\frac{1}{T}\int_{0}^{T}\frac{1}{1-t}\mathrm{d}t\right)^{\frac{1}{2}}$
$\lesssim\log\mathcal{R}\cdot\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{z}}\|\mathbf{s}_{1}-\mathbf{s}_{2}\|^{2}\mathrm{d}t\right)^{\frac{1}{2}}$
$\lesssim\log\mathcal{R}\cdot\|\mathbf{s}_{1}-\mathbf{s}_{2}\|_{L^{\infty}([0,T]\times\mathbb{R}^{d^{*}})}$
$\lesssim\log\mathcal{R}\cdot\|\mathbf{s}_{1}-\mathbf{s}_{2}\|_{L^{\infty}([0,1-{\mathcal{R}}^{-C_{T}}]\times[-\mathcal{O}(1)\sqrt{\log\mathcal{R}},\mathcal{O}(1)\sqrt{\log\mathcal{R}}]^{d^{*}})}.$
@@
$\log\mathcal{N}\left(\mathcal{C},\delta,\|\cdot\|_{L^{\infty}([0,T]\times\mathbb{R}^{d^{*}}}\right)$ $=\log\mathcal{N}\left(\mathcal{C},\delta,\|\cdot\|_{L^{\infty}([0,1-{\mathcal{R}}^{-C_{T}}]\times[-\mathcal{O}(1)\sqrt{\log\mathcal{R}},\mathcal{O}(1)\sqrt{\log\mathcal{R}}]^{d^{*}})}\right)$
$\lesssim JL\log{\left(\frac{LM\kappa(1-{\mathcal{R}}^{-C_{T}}\vee\mathcal{O}(1)\sqrt{\log\mathcal{R}})}{\delta}\right)}$
$\lesssim{\mathcal{R}}^{d^{*}}\log^{13}\mathcal{R}\left(\log^{4}\mathcal{R}+\log\frac{1}{\delta}\right),$
where we used $L=\mathcal{O}(\log^{4}\mathcal{R}),M=\mathcal{O}({\mathcal{R}}^{d^{*}}\log^{7}\mathcal{R}),J=\mathcal{O}(\mathcal{R}^{d^{*}}\log^{9}\mathcal{R}),\kappa=\exp\left(\mathcal{O}(\log^{4}\mathcal{R})\right)$. The details of this derivation can be found in [CJLZ22, Lemma 5.3].
Proof of Lemma 4.4.
Let $\ell(\mathbf{s},\widehat{\mathbf{x}})=\ell_{\mathbf{s}}(\widehat{\mathbf{x}})-\ell_{\mathbf{s}^{*}}(\widehat{\mathbf{x}})$ and $\widehat{\mathcal{X}}^{\prime}=\{\widehat{\mathbf{x}}_{1}^{\prime},\cdots,\widehat{\mathbf{x}}_{n}^{\prime}\}$ be an independent copy of $\widehat{\mathcal{X}}$, then for any $\mathbf{s}_{1}$, $\mathbf{s}_{2}\in\mathcal{C}$, there exists a constant $C_{8}>0$ such that
$|\ell(\mathbf{s}_{1},\widehat{\mathbf{x}})-\ell(\mathbf{s}_{2},\widehat{\mathbf{x})}|=|\ell_{\mathbf{s}_{1}}(\widehat{\mathbf{x}})-\ell_{\mathbf{s}_{2}}(\widehat{\mathbf{x}})|\leq C_{8}\log\mathcal{R}\|\mathbf{s}_{1}-\mathbf{s}_{2}\|_{L^{\infty}([0,T]\times\mathbb{R}^{d^{*}})}.$
@@
where $G(\widehat{\mathbf{s}},\widehat{\mathbf{x}})=\mathbb{E}_{\widehat{\mathcal{X}}^{\prime}}\left[\ell(\widehat{\mathbf{s}},\widehat{\mathbf{x}}^{\prime})-2\ell(\widehat{\mathbf{s}},\widehat{\mathbf{x}})\right].$
Let $\mathcal{C}_{\delta}$ be the $\delta$-covering of $\mathcal{C}$ with minimum cardinality $\mathcal{N}_{\delta}:=\mathcal{N}(\mathrm{NN},\delta,\|\cdot\|_{L^{\infty}([0,T]\times\mathbb{R}^{d^{*}})})$, then for any $\mathbf{s}\in\mathcal{C}$, there exists a $\mathbf{s}_{\delta}\in\mathcal{C}_{\delta}$ such that
$|\ell(\mathbf{s},\widehat{\mathbf{x}})-\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}})|\leq C_{8}\log\mathcal{R}\|\mathbf{s}-\mathbf{s}_{\delta}\|_{L^{\infty}([0,T]\times\mathbb{R}^{d^{*}})}\leq C_{8}\delta\log\mathcal{R}.$
@@
$G(\widehat{\mathbf{s}},\widehat{\mathbf{x}})\leq G(\mathbf{s}_{\delta},\widehat{\mathbf{x}})+3C_{8}\delta\log\mathcal{R}.$
Since $|\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})|\lesssim\log^{2}\mathcal{R}$, there exists a constant $C_{9}>0$ such that $|\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})|\leq C_{9}\log^{2}\mathcal{R}$. We have that $|\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})-\mathbb{E}\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})|\leq 2C_{9}\log^{2}\mathcal{R}$. Let $V^{2}=\mathrm{Var}[\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})]$, then
@@
$\leq C_{8}^{2}\log^{2}\mathcal{R}\mathbb{E}_{\widehat{\mathcal{X}}}\left[\ell_{\mathbf{s}_{\delta}}(\widehat{\mathbf{x}}_{i})-\ell_{\mathbf{s}^{*}}(\widehat{\mathbf{x}}_{i})\right]$
$=C_{8}^{2}\log^{2}\mathcal{R}\mathbb{E}_{\widehat{\mathcal{X}}}[\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})].$
@@
$\mathbb{E}_{\widehat{\mathcal{X}}}[\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})]\geq\frac{V^{2}}{C_{8}^{2}\log^{2}\mathcal{R}}.$
@@
$\leq$ $\leavevmode\nobreak\ \mathbb{P}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\mathbb{E}_{\widehat{\mathcal{X}}}\left[\frac{1}{n}\sum_{i=1}^{n}\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})\right]-\frac{1}{n}\sum_{i=1}^{n}\ell(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})>\frac{t}{2}+\frac{V^{2}}{2C_{8}^{2}\log^{2}\mathcal{R}}\right)$
$\leq$ $\leavevmode\nobreak\ \exp\left(-\frac{nu^{2}}{2V^{2}+\frac{4uC_{9}\log^{2}\mathcal{R}}{3}}\right)$
$\leq$ $\exp\left(-\frac{nt}{8\left(C_{8}^{2}+\frac{C_{9}}{3}\right)\log^{2}\mathcal{R}}\right),$
where $u=\frac{t}{2}+\frac{V^{2}}{2C_{8}^{2}\log^{2}\mathcal{R}}$, and we use $u\geq\frac{t}{2}$ and $V^{2}\leq 2uC_{8}^{2}\log^{2}\mathcal{R}$. Hence, for any $t>3C_{8}\delta\log\mathcal{R}$, we have
$\mathbb{P}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\frac{1}{n}\sum_{i=1}^{n}G(\widehat{\mathbf{s}},\widehat{\mathbf{x}}_{i})>t\right]$ $\leq\mathbb{P}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\mathop{\mathrm{sup}}_{\mathbf{s}\in\mathcal{C}}\frac{1}{n}\sum_{i=1}^{n}G(\mathbf{s},\widehat{\mathbf{x}}_{i})>t\right]$
$\leq\mathbb{P}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\mathop{\mathrm{max}}_{\mathbf{s}_{\delta}\in\mathcal{C}_{\delta}}\frac{1}{n}\sum_{i=1}^{n}G(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})>t-3C_{8}\delta\log\mathcal{R}\right]$
$\leq\mathcal{N}_{\delta}\mathop{\mathrm{max}}_{\mathbf{s}_{\delta}\in\mathcal{C}_{\delta}}\mathbb{P}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\frac{1}{n}\sum_{i=1}^{n}G(\mathbf{s}_{\delta},\widehat{\mathbf{x}}_{i})>t-3C_{8}\delta\log\mathcal{R}\right]$
$\leq\mathcal{N}_{\delta}\exp\left(-\frac{n(t-3C_{8}\delta\log\mathcal{R})}{8\left(C_{8}^{2}+\frac{C_{9}}{3}\right)\log^{2}\mathcal{R}}\right).$
By setting $a=\left[3C_{8}\log\mathcal{R}+8\left(C_{8}^{2}+\frac{C_{9}}{3}\right)\log^{2}\mathcal{R}\log\mathcal{N}_{\delta}\right]\delta$ and $\delta=\frac{1}{n}$, then we obtain
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\frac{1}{n}\sum_{i=1}^{n}G(\widehat{\mathbf{s}},\widehat{\mathbf{x}}_{i})\right]$ $\leq a+\mathcal{N}_{\delta}\int_{a}^{\infty}\exp\left(-\frac{n(t-3C_{8}\delta\log\mathcal{R})}{8\left(C_{8}^{2}+\frac{C_{9}}{3}\right)\log^{2}\mathcal{R}}\right)\mathrm{d}t$ (21)
$\leq\left[3C_{8}\log\mathcal{R}+8\left(C_{8}^{2}+\frac{C_{9}}{3}\right)\log^{2}\mathcal{R}\log\mathcal{N}_{\delta}\right]\delta+\frac{8\left(C_{8}^{2}+\frac{C_{9}}{3}\right)\log^{2}\mathcal{R}}{n}$
$\lesssim{\mathcal{R}}^{d^{*}}\log^{15}\mathcal{R}\left(\log^{4}\mathcal{R}+\log\frac{1}{\delta}\right)\delta+\frac{\log^{2}\mathcal{R}}{n}$
$\lesssim\frac{{\mathcal{R}}^{d^{*}}\log^{15}\mathcal{R}\left(\log^{4}\mathcal{R}+\log n\right)}{n}.$
@@
$\ell_{\widehat{\mathbf{s}}}^{\mathrm{trunc}}(\widehat{\mathbf{x}}_{i})=\mathbb{E}_{\mathbf{z}}\left(\frac{1}{T}\int_{0}^{T}\left\|\widehat{\mathbf{s}}(1-t,\widehat{\mathbf{x}}_{i}+\sigma\sqrt{1-t}\mathbf{z})+\frac{\mathbf{z}}{\sigma\sqrt{1-t}}\right\|^{2}\mathrm{d}t\cdot\mathbf{I}_{\{\|\mathbf{z}\|_{\infty}\leq r\}}\right),$
@@
^$\widehat{\ell}_{\widehat{\mathbf{s}}}^{\mathrm{trunc}}(\widehat{\mathbf{x}}_{i})=\frac{1}{m}\sum_{j=1}^{m}\left\|\widehat{\mathbf{s}}(1-t_{j},\widehat{\mathbf{x}}_{i}+\sigma\sqrt{1-t_{j}}\mathbf{z}_{j})+\frac{\mathbf{z}_{j}}{\sigma\sqrt{1-t_{j}}}\right\|^{2}\mathbf{I}_{\{\|\mathbf{z}_{j}\|_{\infty}\leq r\}}.$
@@
$(A)$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\mathbf{z}}\left(\frac{1}{T}\int_{0}^{T}\left\|\widehat{\mathbf{s}}(1-t,\widehat{\mathbf{x}}_{i}+\sigma\sqrt{1-t}\mathbf{z})+\frac{\mathbf{z}}{\sigma\sqrt{1-t}}\right\|^{2}\mathrm{d}t\cdot\mathbf{I}_{\{\|\mathbf{z}\|_{\infty}>r\}}\right)$
$\lesssim\left(\log\mathcal{R}\mathbb{P}(\|\mathbf{z}\|_{\infty}>r)+\mathbb{E}\left(\|\mathbf{z}\|^{2}\mathbf{I}_{\{\|\mathbf{z}\|_{\infty}>r\}}\right)\right)\cdot\frac{1}{T}\int_{0}^{T}\frac{1}{\sigma^{2}(1-t)}\mathrm{d}t$
$\lesssim\log\mathcal{R}\left(\log\mathcal{R}\mathbb{P}(\|\mathbf{z}\|_{\infty}>r)+[\mathbb{E}(\|\mathbf{z}\|^{4})]^{\frac{1}{2}}\mathbb{P}({\|\mathbf{z}\|_{\infty}>r})^{\frac{1}{2}}\right)$
$\lesssim\log^{2}\mathcal{R}\mathbb{P}({\|\mathbf{z}\|_{\infty}>r})^{\frac{1}{2}}$
$\lesssim\log^{2}\mathcal{R}\exp\left(-\frac{r^{2}}{4}\right).$
Therefore, there exists a constant $C_{10}>0$ such that
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\frac{1}{n}\sum_{i=1}^{n}\left(\ell_{\widehat{\mathbf{s}}}(\widehat{\mathbf{x}}_{i})-\ell_{\widehat{\mathbf{s}}}^{\mathrm{trunc}}(\widehat{\mathbf{x}}_{i})\right)\right)\leq C_{10}\log^{2}\mathcal{R}\exp\left(-\frac{r^{2}}{4}\right).$
Let $h_{\mathbf{s}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})=\|\mathbf{s}(1-t,\widehat{\mathbf{x}}_{i}+\sigma\sqrt{1-t}\mathbf{z})+\frac{\mathbf{z}}{\sigma\sqrt{1-t}}\|^{2}\mathbf{I}_{\{\|\mathbf{z}\|_{\infty}\leq r\}}$, then there exists a constant $C_{11}>0$ such that
$0\leq h_{\mathbf{s}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})\lesssim\frac{r^{2}+\log\mathcal{R}}{1-T}\leq C_{11}{\mathcal{R}}^{C_{T}}(r^{2}+\log\mathcal{R}):=E_{\mathcal{R}}(r).$
For any $\delta_{1}>0$, $\mathbf{s}\in\mathrm{NN}$, there exists a $\mathbf{s}_{\delta_{1}}\in\mathcal{C}_{\delta_{1}}$ and a constant $C_{12}>0$ such that
$|h_{\mathbf{s}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})-h_{\mathbf{s}_{\delta_{1}}}(t,\widehat{\mathbf{x}}_{i},\mathbf{z})|$ $\lesssim\delta\left\|\mathbf{s}+\mathbf{s}_{\delta_{1}}+\frac{2\mathbf{z}}{\sigma\sqrt{1-t}}\right\|\mathbf{I}_{\{\|\mathbf{z}\|_{\infty}\leq r\}}$
$\lesssim\frac{r+\sqrt{\log\mathcal{R}}}{\sqrt{1-T}}\cdot\delta_{1}$
$\leq C_{12}{\mathcal{R}}^{\frac{C_{T}}{2}}(r+\sqrt{\log\mathcal{R}})\delta_{1}.$